OpenStudy (anonymous):

1. The distance traveled by an object can be modeled by the equation d = ut + 0.5at2 where d = distance, u = initial velocity, t = time, and a = acceleration. Solve this formula for a. Show all steps in your work.

OpenStudy (cwrw238):

first isolate the term containing a subtracting ut from either side will do this:- d - ut = ut - ut +0.5at^2 d - ut = 0.5at^2 now divide both sides by 0.5t^2 and you have it

OpenStudy (anonymous):



OpenStudy (anonymous):

not s its d

OpenStudy (cwrw238):

@Sambhavvinaykya - please don;t just give the answer - plz read the sites Code of Conduct

OpenStudy (anonymous):

Thanks guys!!

OpenStudy (anonymous):

ohh sorry i will make sure i explain it next time

OpenStudy (cwrw238):

yw